x^2+xy=5, y^2+xy=3
Answer by dkppathak(439)

x^2+xy=5,(i)
y^2+xy=3(ii)
taking x and y common from i and ii
x(x+y)=5 y(y+x)=3
x+y=5/x(iii) and x+y=3/y(iv)
from these equations
5/x=3/y
y=3x/5
putting the value of y in equation (iii)
x(x+3x/5)=5
x(8x/5)=5
8x^2=25
x^2=25/8
x=sqrt 25/8
simlarly
we can find x=5y/3
putting value in (iv)
y(y+5y/3)=3
8y^2=9
y^2=9/8
y=sqrt 9/8
answer x=sqrt25/8
y=sqrt 9/8